OpenStudy (anonymous):

x^2 - (root2+1)x+ root2=0 solve by completing the square.

OpenStudy (anonymous):

\[x ^{2}-(\sqrt{2}+1)x+\sqrt{2}=0\]

OpenStudy (anonymous):

You want to make it fit the form \[(x+p)^{2} = x^{2}+2xp+p^{2}\] by definition of completing the square, yah? You can rewrite your middle term as \[+(-\sqrt{2}-1)x\] to see what p is defined as more clearly. Does that help?

OpenStudy (anonymous):

so \[2p = (-\sqrt{2}-1)\], then just plug and chug p^2 to figure out what the right hand side should look like.